# GraphAllocBench: A Flexible Benchmark for Preference-Conditioned Multi-Objective Policy Learning

### v2

@@ 2.1 Pareto Front
In multi-objective optimization, a solution is Pareto optimal if it cannot be improved in any objective without degrading at least one other objective. Let’s consider a multi-objective maximization problem with $N$ objectives, represented by a vector function:
@@ 3.1 Definitions
Assignment of resources to demands, represented by allocation matrix $A$, where $A_{i,j}$ is the amount of $R_{i}$ given to $D_{j}$. The observation returned by the environment is the normalized allocation matrix, concatenated with the current preferences vector for each rollout.
@@
Multi-objective reward signal at the end of each step (i.e. after each action), defined by $\mathbf{J}(\mathbf{P}):\mathbb{R}^{|\mathbf{P}|}\to\mathbb{R}^{N}$, mapping $|\mathbf{P}|=|D|$ productions to $N$ objective values.
@@ 4 GraphAllocBench
For each sampled preference, we roll out the agent’s policy deterministically for $T$ steps to obtain the final allocation of resources to demands. The Das and Dennis method  Das and Dennis (1998) generates a set of evenly distributed preferences by constructing a structured grid of points on an $N$-dimensional simplex (for $N$ objectives). Hence, the number of preferences sampled increases combinatorially with $N$. For scenarios where it is computationally feasible to compute the allocations that lead to solutions on the analytical Pareto Front, we compare the hypervolume of the set of predicted solutions for the sampled preferences with the ideal Pareto Front. The evaluation metrics are defined respectively in the following sections.
@@ 4.2 Proportion of Non-Dominated Solutions (PNDS)
While this metric is beneficial as a supplement to the hypervolume, it still struggles when the number of objectives is high due to sampling inefficiency and requires sampling exponentially more preferences using  Das and Dennis (1998) with increasing $N$ to ensure fair comparisons.
@@ 5.1 Main Problem Set
A PCPL agent may find it difficult to deal with sparse rewards, especially when optimizing a continuous preference vector. An example of this would be the $\mathbf{J}_{i}(\mathbf{P})=\text{max}(0,\mathbf{P}_{j}-5)$ floor function, where we receive no reward for Objective $i$ until we produce more than $5$ units for Demand $j$. Instead of deciding whether to a) optimize a particular objective by creating productions until the reward spikes or to b) allocate no resources to the given production, the agent might create insufficient production, such that resources are wasted but no additional reward is received. This results in a lower proportion of non-dominated solutions, as observed for Problem 1c in Figure 2. In Figure 3, we also observe many predicted solutions along the horizontal axis where the reward component for Objective 1 is 0. However, there should only be one Pareto-optimal solution along the horizontal axis that maximizes Objective 0.
@@ Appendix B Ordering Score Computation
1:Environment envenv with $N$ objectives, policy $\pi_{\theta}$, number of samples $n_{\text{samp}}$, steps $n_{\text{step}}$ in each parameter sweep

### v3

@@ 2.1 Pareto Front
In multi-objective optimization, a solution is Pareto optimal if it cannot be improved in any objective without degrading at least one other objective. Let’s consider a multi-objective maximization problem with NN objectives, represented by a vector function:
@@ 3.1 Definitions
Assignment of resources to demands, represented by allocation matrix AA, where $A_{i,j}$ is the amount of $R_{i}$ given to $D_{j}$. The observation returned by the environment is the normalized allocation matrix, concatenated with the current preferences vector for each rollout.
@@
Multi-objective reward signal at the end of each step (i.e. after each action), defined by $\mathbf{J}(\mathbf{P}):\mathbb{R}^{|\mathbf{P}|}\to\mathbb{R}^{N}$, mapping $|\mathbf{P}|=|D|$ productions to NN objective values.
@@ 4 GraphAllocBench
For each sampled preference, we roll out the agent’s policy deterministically for $T$ steps to obtain the final allocation of resources to demands. The Das and Dennis method  Das and Dennis (1998) generates a set of evenly distributed preferences by constructing a structured grid of points on an NN-dimensional simplex (for NN objectives). Hence, the number of preferences sampled increases combinatorially with NN. For scenarios where it is computationally feasible to compute the allocations that lead to solutions on the analytical Pareto Front, we compare the hypervolume of the set of predicted solutions for the sampled preferences with the ideal Pareto Front. The evaluation metrics are defined respectively in the following sections.
@@ 4.2 Proportion of Non-Dominated Solutions (PNDS)
While this metric is beneficial as a supplement to the hypervolume, it still struggles when the number of objectives is high due to sampling inefficiency and requires sampling exponentially more preferences using  Das and Dennis (1998) with increasing NN to ensure fair comparisons.
@@ 5.1 Main Problem Set
A PCPL agent may find it difficult to deal with sparse rewards, especially when optimizing a continuous preference vector. An example of this would be the $\mathbf{J}_{i}(\mathbf{P})=\text{max}(0,\mathbf{P}_{j}-5)$ floor function, where we receive no reward for Objective $i$ until we produce more than 55 units for Demand $j$. Instead of deciding whether to a) optimize a particular objective by creating productions until the reward spikes or to b) allocate no resources to the given production, the agent might create insufficient production, such that resources are wasted but no additional reward is received. This results in a lower proportion of non-dominated solutions, as observed for Problem 1c in Figure 2. In Figure 3, we also observe many predicted solutions along the horizontal axis where the reward component for Objective 1 is 0. However, there should only be one Pareto-optimal solution along the horizontal axis that maximizes Objective 0.
@@ Appendix B Ordering Score Computation
1:Environment envenv with NN objectives, policy $\pi_{\theta}$, number of samples $n_{\text{samp}}$, steps $n_{\text{step}}$ in each parameter sweep